Shift report 20011231 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Dec 31 2001 - 16:06:51 EST

  • Next message: Flemming Videbaek: "Fw: Happy New Year!! from scheduling physicist"

    ------------------------------------------------------------
    Date: Mon Dec 31 16:06:51 2001, Shift: 08:00-16:00
    Supervisor: sjs, with: 
    Log book#11 pages 126 - 127
    
    Shift Summary:
    ------------------------------------------------------------
    Stable beam until 12:30, then MCR called saying they wanted
    to dump ASAP. (DB called for explanation and told 
    MCR wanted to allow ring access for experiments before
    New Years Eve...)
    
    No beam for remainder of shift.
    
    Run 6220. 32.5K trig. 6 FS at 12 deg, 1/8A, MRS at 90 deg.
    Run 6221  no change in config. 16.5K trig. 6
    
    Dana placed inner elastic counters into logic to understand
    difference in BRAHMS and pp2pp rates, but
    left trigger as initially found.
    
    No new detector problems uncovered.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Mon Dec 31 2001 - 16:07:27 EST